OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4, 2).

OpenStudy (anonymous):

I wonder whether you're more upset that she's not doing work or that she's not giving medals... ;p

OpenStudy (amistre64):

lol ..... medals are a dime a dozen, but there are others that answer her as well with no 'rewards'. the thing im upset about is her demeanor when asked to work thru the problem.

OpenStudy (amistre64):

the "shutup and answer it" mentality is a bit annoying.

OpenStudy (anonymous):

yes, I agree. I wish I had an ignore feature so I could choose to not be distracted by such people who aren't really interested in learning.

OpenStudy (amistre64):

:) delaine thnx for the medal :) now, to answer the problem.... we just have to apply the information given into the formula for the line...

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

Do you know the equation for the point slope formula?

OpenStudy (amistre64):

y = mx+b is the slope -intercept form; and given the slope and a point we can just determine its equations by this: \[y = m(x-x_o)+y_o\] y = 2(x-4)+2 y = 2x -8+2 y = 2x -6

OpenStudy (anonymous):

y=mx+b right?

OpenStudy (anonymous):

no that's slope intercept.

OpenStudy (anonymous):

ist that what you adked?

OpenStudy (anonymous):

point-slope is: \[y - y_1 = m(x - x_1)\] Where \(x_1,y_1\) are the x,y values of a point on the line, and m is the slope.

OpenStudy (anonymous):

*asked

OpenStudy (anonymous):

So in your case you have (4,2) is the point and 2 is the slope, so: y - 2 = 2(x - 4) Then solve for y (and simplify) and you should have slope intercept.

OpenStudy (amistre64):

the use of \(y_1\) and \(x_1\) are common, but not crucial 'forms'; i used \(y_o\) and \(x_0\) to name mine.